OpenStudy (nessapetrenko):

Solve. 3(x + 1) - 2x = -6. A) x = 1 B) x = 5 Eliminate C) x = -7 D) x = -9

OpenStudy (maddy1251):

First, why don't we get rid of them parentheses. That will make it cleaner looking, right? :)

OpenStudy (maddy1251):

3(x + 1) - 2x = -6. 3x+3-2x=-6 Woo, looks A LOT better now. So, we want to solve for 'x'. We want 'x' by itself. Lets work on that, shall we?

OpenStudy (maddy1251):

3x+3-2x=-6 Lets get that +3 away. 3x+3-3-2x=-6+3 3x-2x=-3

OpenStudy (maddy1251):

Now, we have like terms on each side, the 'x' isn't mixed up with any numbers that do not have the 'x' terms. Lets continue with 3x-2x=-3 So now we will subtract the 3x-2x because they are like terms.. That will yield '1x' or simply 'x' So x=-3

OpenStudy (maddy1251):

Hmm, but that isn't an answer..

OpenStudy (maddy1251):

Well.. Since that didn't work, we will plug-and-chug.

OpenStudy (nessapetrenko):

okay!

OpenStudy (maddy1251):

The answer is D, I guess. Just plug in what x=n, into the original equation. I shall show you

OpenStudy (nessapetrenko):

k thx.

OpenStudy (maddy1251):

\[3(x + 1) - 2x = -6. : 3(-9+1)-2(-9)=-6\]

OpenStudy (maddy1251):

Sorry for the junk before. Not sure why it didn't yield the answer before.